# Conditional Extreme Value Estimation for Dependent Time Series

### v2

@@ 1 Introduction
Functional limit theorems for the tail empirical process under dependence conditions have been studied in the literature. Early results were established by Rootzen_1995 for $\alpha$-mixing sequences and later extended by Rootzen_2009 to more general dependence structures. In Drees_2000 the asymptotic normality of the POT estimator was shown under $\beta$-mixing conditions. Among the most widely used estimators for heavy-tailed data is the Hill estimator, originally introduced by Hill1975. Its consistency in dependent settings was first investigated by Resnick_Starica_1995, who analysed its behaviour under various dependence structures. Further work by Resnick_Starica_1997 established the consistency of the tail empirical measure for autoregressive processes with heavy-tailed innovations. The asymptotic normality of the Hill estimator for i.i.d. data has been well documented, with significant contributions from Hall_1982Davis_Resnick_1984Haeusler_Teugels_1985, and Csorgo_Mason_1985. Extending these results to stationary sequences, HSING1991117 proved a central limit theorem for the Hill estimator under $\beta$-mixing conditions, and a modern treatment can be found in Kulik2020.
@@
Other work is related to ours. In the presence of covariates, the asymptotic properties of extreme value estimators have been investigated using regression-based approaches. In Daouia2010 tail index estimation is studied under $\alpha$-mixing conditions using a smoothed Pickands estimator, a technique later generalized by Daouia_2013; Daouia_Stupfler_Usseglio-Carleve_2023 to regression models within the general max-domain of attraction. Furthermore, tail index regression in a random covariate setting has been analysed using a maximum likelihood framework, as explored by Wang01092009. Finally, asymptotic normality of the conditional kernel Hill estimator has been explored through different techniques and for a special second-order regularly varying class in the preprint GoegebeurGuillou2024.
@@ 3.2 Central limit theorems
Notice that the speed of convergence is determined by the effective sample size $nh_{n}\overline{F}^{x}(u_{n}^{x})=h_{n}k_{n}$, which is slower than in the unconditional case, which is simply $k_{n}$. This price to pay, however, results in a variance which agrees with the i.i.d. conditional case, in contrast to the more involved expression for unconditional time series provided as an infinite sum, e.g. in Kulik2020.
@@ 4.2 Asymptotic normality
In essence, Lemmas 1 and 2 transfer the asymptotic bias conditions into additional speed conditions on the $h_{n}$ and $k_{n}$ sequences. Standard second-order conditions in extremes are often formulated in terms of limits which imply (4); in particular the broad Hall class (HallWelsh1985) falls into this framework. Second-order conditions akin to (2) are not often encountered in the literature, and our result is inspired by (Kulik2020, Eq. (9.6.3)). The result of Lemma 3 provides an alternative way of verifying the anti-clustering conditions. For convenience, we now restate the main speed conditions for asymptotic normality of the conditional Hill estimator to hold.
@@
It resembles the rate achieved in the unconditional case under a corresponding second order assumption, $k_{n}=o(n^{\rho})$ where $\rho=2(\zeta_{x}-1/\gamma(x))/(2\zeta_{x}-1/\gamma(x))$ (see Kulik2020, problem 9.6), however the fraction $2/3$ in the exponent is caused by the bandwidth sequence slowing the speed of convergence.
The preprint GoegebeurGuillou2024 obtains similar asymptotic rates for the conditional Hill estimator. We now compare their main assumptions. Their kernel assumptions $(\mathcal{K})$ are similar to ours; they also cover the multidimensional covariate case. They throughout assume a second order condition $(\mathcal{D})$, which is seen to be a special case of equation (4); they do not cover the second order condition of equation (6). Their assumption $(\mathcal{J})$ implies our anti-clustering condition 2. Their strong condition $(\mathcal{X})$ on the joint behaviour of $(X_{0},X_{i})$ is here replaced by simply $g_{X_{0},X_{i}}$ being bounded for every $i$. Their condition (A) is similar to our beta-mixing condition 3 but it depends on an involved probability sequence $\{\nu_{n}\}$. Their condition (B) implies our anti-clustering condition 6. In their Theorem 3.2 and 3.3, the first bandwidth condition corresponds to condition 2, the second to equation (16), the third and fourth to the first and second part of equation (5).
@@ 5.1 Bandwidth selection
    Sheather-Jones bandwidth fitted only on the covariate process (see Sheather_Jones). It’s a plug-in method that provides a global bandwidth, independent of $k_{n}$.
@@
    Various approaches concerning non-parametric regression (see Daouia2010; Gannoun2002 ) have proposed a purely data-driven, cross-validation bandwidth selection method
@@
    Alternatively, as proposed by Li_Racine, a leave-one-out doubly-smoothed Nadaraya–Watson estimator of the conditional survival function, where also $\left\{Y_{j}\right\}$ is smoothed with distribution function $G$, using bandwidth $b_{n}$ is given by
@@ 5.2 Simulation study
The conditional Fréchet and Pareto are known to satisfy the second-order condition of equation (4). By the remarks on page 224 in Mikosch_Wintenberger, it holds that $V^{(i)}_{j}:=e^{W_{j}^{(i)}-(1/\gamma(X_{j}))\sigma_{j}^{2}/2}$ in equation (12) can be represented by $(f_{j}(U_{i}))_{j=1}^{\infty}$ for suitable non-negative measurable functions $f_{j}$ on $(0,1)$, and $\left\{U_{i}\right\}$ are i.i.d. uniform variables on $(0,1)$. By eq. (5.5.7) therein, the CSGMS process above has conditional Fréchet marginals. In particular it also satisfies the condition of equation (4) and we note that all the three constructions satisfy equation (1). The mapping $z\mapsto 1/\gamma(z)=1/(3z(z-1)+1)$ is clearly Lipschitz continuous and hence condition 1 holds. Condition 6 and equation (8) are verified for the conditional Fréchet and Pareto in Lemma 9, Appendix C, and anti-clustering conditions for the CSGMS hold due to Lemma 10, Appendix C. In the first two cases, the underlying processes are simulated from $\text{AR}(1)$-processes with auto-regressive coefficients absolutely bounded by $1$, which implies $\beta$-mixing with mixing coefficients of exponential decay (see Ibragimov1962). When considering the unconditional version of the CSGMS-process, it is possible to show $\beta$-mixing decay of exponential order (see Kulik2020, Example 13.5.4). In equation (12) we extend the construction by bundling with the covariate sequence, which itself is i.i.d and in particular of exponential $\beta$-mixing decay. Since $z\mapsto 1/\gamma(z)$ is Lipschitz, then also the bivariate sequence $\left\{X_{j},Y_{j}\right\}$ is of exponential $\beta$-mixing decay. Hence, since the conditional slowly varying functions in equation (1) are seen to satisfy condition 3 for both Pareto and Fréchet marginals, we may invoke Theorem 8 to conclude that there exists sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n})\right\}$ such that all conditions required in Theorem 5 are met in the above simulation studies.
@@ 6.1 Description of dataset
The West Texas Intermediate (WTI) is a grade of crude oil and is considered one of the main benchmarks in oil pricing. We are interested to examine whether the spot price of WTI is conditionally regularly varying given the Crude Oil Volatility Index (CBOE OVX) and, if this is the case, how the tail index is influenced by different levels of implied volatility. The CBOE OVX is constructed by applying the VIX (Chicago Board Options Exchange’s Volatility Index) methodology (see whaley2000investor) to options on the United States Oil Fund (USO), using a weighted average of mid-quote prices from calls and puts across a range of strike prices. This aggregated implied volatility is extrapolated to a 30‐day horizon to provide a forward-looking measure of the expectation of the market of price fluctuations and is expressed in percent.
Now, as USO is designed to track the WTI price using future contracts, the OVX provides a forward looking market expectation of the WTI price volatility. Thus, in the case of regularly varying marginal distributions of the WTI, we expect that the tail index varies as a function of OVX. While the OVX is mean-reverting and can be considered stationary (see chen2018predictive), oil prices are generally not stationary, and are consequently often transformed to log-returns (see hamilton1983oil; sadorsky1999oil) to achieve stationarity. We follow this approach. The corresponding time series are depicted in Figure 4. Note how high activity periods such as the financial crisis (2008), oil price collapse due to US and OPEC oil war (2015), covid period (2020-2021), and the Russian and Ukrainian conflict (2022), are visually observable. Remark that the WTI price attained a negative value on the 20th of April 2020, however this observation has been removed in order to transform with logarithms.
@@ 6.2 Analysis of crude oil spot price
In Figure 6 we consider the conditional Hill estimator of the WTI log-returns given the uniformly transformed OVX. We observe that the largest $\widehat{\gamma}_{n}(x)$ value is attained for both negative and positive sides of the WTI log-returns, when the implied volatility is large. This is in accordance with the common financial intuition that large implied volatility implies heavier tails (and thus fewer moments) for log-returns of markets on which it is based. Furthermore, when the implied volatility is large, the negative side of the WTI log-returns exhibits slightly heavier tails than the positive side. This has been documented for many assets (see Christie_82), where shock effects can create large upswings but even greater downswings. In addition, the confidence intervals that arise from Theorem 5 are depicted. For the both sides, the lower bound of the estimator corresponding to high volatility is separated from the upper bound of the of the estimators, corresponding to low and mid volatility, within the stable regions. This supports the claim the the the WTI log-returns change tail index according the implied volatility. Since the conditional Hill estimator can be viewed as a function of the covariate value $x$ for a fixed sample fraction $k_{n}$ bivariately, we also plot these “risk profiles”, where slightly noisy right-skewed smiles can be seen, which indicates that the tail index is large when volatility is low and even larger when high.
@@ 7 Conclusion and outlook
Apart from considering stronger intertemporal dependence structures, our work also allows us to consider extensions in several directions. First, having an estimate of the entire tail at our disposal our results can be used to construct and statistically analyse conditional risk measures such as conditional expected shortfall or conditional tail moments. Secondly, a promising direction is to consider specific models where the same time series can be used as the covariate and response, but at different lags; Markov processes are particularly attractive in this connection. Finally, removing top observations can sometimes lead to the construction of stable estimators that allow transparent threshold selection as in bladt2020threshold, which is yet to be explored both for univariate and conditional cases.
@@ Appendix B Proofs of results from Section 4
using Bingham1987 Proposition 1.3.6 (i). Hence $\log\left(n/k_{n}\right)\sim(1/\gamma(x))\log(u_{n}^{x})$ and consequently
@@ Appendix E Background on fundamental limit theorems
We introduce $\alpha$- and $\beta$-mixing using Rosenblatt1956ACL; Volkonskii1959 as primary references (see also Bradley_05; Kulik2020) and state useful results on $\beta$-mixing sequences.
@@
In this section we present existing results on regarding Central Limit Theorems using BillingsleyPatrick2009Copm and vanderVaart2023 as primary reference, see also Kulik2020.

### v3

@@ 1 Introduction
Functional limit theorems for the tail empirical process under dependence conditions have been studied in the literature. Early results were established by Rootzén (1995) for $\alpha$-mixing sequences and later extended by Rootzén (2009) to more general dependence structures. In Drees (2000) the asymptotic normality of the POT estimator was shown under $\beta$-mixing conditions. Among the most widely used estimators for heavy-tailed data is the Hill estimator, originally introduced by Hill (1975). Its consistency in dependent settings was first investigated by Resnick and Stărică (1995), who analysed its behaviour under various dependence structures. Further work by Resnick and Stărică (1997) established the consistency of the tail empirical measure for autoregressive processes with heavy-tailed innovations. The asymptotic normality of the Hill estimator for i.i.d. data has been well documented, with significant contributions from Hall (1982)Davis and Resnick (1984)Haeusler and Teugels (1985), and Csörgő and Mason (1985). Extending these results to stationary sequences, Hsing (1991) proved a central limit theorem for the Hill estimator under $\beta$-mixing conditions, and a modern treatment can be found in Kulik and Soulier (2020).
@@
Other work is related to ours. In the presence of covariates, the asymptotic properties of extreme value estimators have been investigated using regression-based approaches. In Daouia et al. (2010) tail index estimation is studied under $\alpha$-mixing conditions using a smoothed Pickands estimator, a technique later generalized by Daouia et al. (2013, 2023) to regression models within the general max-domain of attraction. Furthermore, tail index regression in a random covariate setting has been analysed using a maximum likelihood framework, as explored by Wang and Tsai (2009). Finally, asymptotic normality of the conditional kernel Hill estimator has been explored through different techniques and for a special second-order regularly varying class in the preprint Goegebeur and Guillou (2024).
@@ 3.2 Central limit theorems
Notice that the speed of convergence is determined by the effective sample size $nh_{n}\overline{F}^{x}(u_{n}^{x})=h_{n}k_{n}$, which is slower than in the unconditional case, which is simply $k_{n}$. This price to pay, however, results in a variance which agrees with the i.i.d. conditional case, in contrast to the more involved expression for unconditional time series provided as an infinite sum, e.g. in Kulik and Soulier (2020, Eq. (9.5.4)).
@@ 4.2 Asymptotic normality
In essence, Lemmas 1 and 2 transfer the asymptotic bias conditions into additional speed conditions on the $h_{n}$ and $k_{n}$ sequences. Standard second-order conditions in extremes are often formulated in terms of limits which imply (4); in particular the broad Hall class (Hall and Welsh, 1985) falls into this framework. Second-order conditions akin to (2) are not often encountered in the literature, and our result is inspired by (Kulik and Soulier, 2020, Eq. (9.6.3)). The result of Lemma 3 provides an alternative way of verifying the anti-clustering conditions. For convenience, we now restate the main speed conditions for asymptotic normality of the conditional Hill estimator to hold.
@@
It resembles the rate achieved in the unconditional case under a corresponding second order assumption, $k_{n}=o(n^{\rho})$ where $\rho=2(\zeta_{x}-1/\gamma(x))/(2\zeta_{x}-1/\gamma(x))$ (see Kulik and Soulier (2020), problem 9.6), however the fraction $2/3$ in the exponent is caused by the bandwidth sequence slowing the speed of convergence.
The preprint Goegebeur and Guillou (2024) obtains similar asymptotic rates for the conditional Hill estimator. We now compare their main assumptions. Their kernel assumptions $(\mathcal{K})$ are similar to ours; they also cover the multidimensional covariate case. They throughout assume a second order condition $(\mathcal{D})$, which is seen to be a special case of equation (4); they do not cover the second order condition of equation (6). Their assumption $(\mathcal{J})$ implies our anti-clustering condition 2. Their strong condition $(\mathcal{X})$ on the joint behaviour of $(X_{0},X_{i})$ is here replaced by simply $g_{X_{0},X_{i}}$ being bounded for every $i$. Their condition (A) is similar to our beta-mixing condition 3 but it depends on an involved probability sequence $\{\nu_{n}\}$. Their condition (B) implies our anti-clustering condition 6. In their Theorem 3.2 and 3.3, the first bandwidth condition corresponds to condition 2, the second to equation (16), the third and fourth to the first and second part of equation (5).
@@ 5.1 Bandwidth selection
    Sheather-Jones bandwidth fitted only on the covariate process (see Sheather and Jones (1991)). It’s a plug-in method that provides a global bandwidth, independent of $k_{n}$.
@@
    Various approaches concerning non-parametric regression (see Daouia et al. (2010); Gannoun et al. (2002) ) have proposed a purely data-driven, cross-validation bandwidth selection method
@@
    Alternatively, as proposed by Li and Racine (2008), a leave-one-out doubly-smoothed Nadaraya–Watson estimator of the conditional survival function, where also $\left\{Y_{j}\right\}$ is smoothed with distribution function $G$, using bandwidth $b_{n}$ is given by
@@ 5.2 Simulation study
The conditional Fréchet and Pareto are known to satisfy the second-order condition of equation (4). By the remarks on page 224 in Mikosch and Wintenberger (2024), it holds that $V^{(i)}_{j}:=e^{W_{j}^{(i)}-(1/\gamma(X_{j}))\sigma_{j}^{2}/2}$ in equation (12) can be represented by $(f_{j}(U_{i}))_{j=1}^{\infty}$ for suitable non-negative measurable functions $f_{j}$ on $(0,1)$, and $\left\{U_{i}\right\}$ are i.i.d. uniform variables on $(0,1)$. By eq. (5.5.7) therein, the CSGMS process above has conditional Fréchet marginals. In particular it also satisfies the condition of equation (4) and we note that all the three constructions satisfy equation (1). The mapping $z\mapsto 1/\gamma(z)=1/(3z(z-1)+1)$ is clearly Lipschitz continuous and hence condition 1 holds. Condition 6 and equation (8) are verified for the conditional Fréchet and Pareto in Lemma 9, Appendix C, and anti-clustering conditions for the CSGMS hold due to Lemma 10, Appendix C. In the first two cases, the underlying processes are simulated from $\text{AR}(1)$-processes with auto-regressive coefficients absolutely bounded by $1$, which implies $\beta$-mixing with mixing coefficients of exponential decay (see Ibragimov (1962)). When considering the unconditional version of the CSGMS-process, it is possible to show $\beta$-mixing decay of exponential order (see Kulik and Soulier (2020), Example 13.5.4). In equation (12) we extend the construction by bundling with the covariate sequence, which itself is i.i.d and in particular of exponential $\beta$-mixing decay. Since $z\mapsto 1/\gamma(z)$ is Lipschitz, then also the bivariate sequence $\left\{X_{j},Y_{j}\right\}$ is of exponential $\beta$-mixing decay. Hence, since the conditional slowly varying functions in equation (1) are seen to satisfy condition 3 for both Pareto and Fréchet marginals, we may invoke Theorem 8 to conclude that there exists sequences $\left\{(k_{n},h_{n},r_{n},\ell_{n})\right\}$ such that all conditions required in Theorem 5 are met in the above simulation studies.
@@ 6.1 Description of dataset
The West Texas Intermediate (WTI) is a grade of crude oil and is considered one of the main benchmarks in oil pricing. We are interested to examine whether the spot price of WTI is conditionally regularly varying given the Crude Oil Volatility Index (CBOE OVX) and, if this is the case, how the tail index is influenced by different levels of implied volatility. The CBOE OVX is constructed by applying the VIX (Chicago Board Options Exchange’s Volatility Index) methodology (see Whaley (2000)) to options on the United States Oil Fund (USO), using a weighted average of mid-quote prices from calls and puts across a range of strike prices. This aggregated implied volatility is extrapolated to a 30‐day horizon to provide a forward-looking measure of the expectation of the market of price fluctuations and is expressed in percent.
Now, as USO is designed to track the WTI price using future contracts, the OVX provides a forward looking market expectation of the WTI price volatility. Thus, in the case of regularly varying marginal distributions of the WTI, we expect that the tail index varies as a function of OVX. While the OVX is mean-reverting and can be considered stationary (see Chen et al. (2018)), oil prices are generally not stationary, and are consequently often transformed to log-returns (see Hamilton (1983); Sadorsky (1999)) to achieve stationarity. We follow this approach. The corresponding time series are depicted in Figure 4. Note how high activity periods such as the financial crisis (2008), oil price collapse due to US and OPEC oil war (2015), covid period (2020-2021), and the Russian and Ukrainian conflict (2022), are visually observable. Remark that the WTI price attained a negative value on the 20th of April 2020, however this observation has been removed in order to transform with logarithms.
@@ 6.2 Analysis of crude oil spot price
In Figure 6 we consider the conditional Hill estimator of the WTI log-returns given the uniformly transformed OVX. We observe that the largest $\widehat{\gamma}_{n}(x)$ value is attained for both negative and positive sides of the WTI log-returns, when the implied volatility is large. This is in accordance with the common financial intuition that large implied volatility implies heavier tails (and thus fewer moments) for log-returns of markets on which it is based. Furthermore, when the implied volatility is large, the negative side of the WTI log-returns exhibits slightly heavier tails than the positive side. This has been documented for many assets (see Christie (1982)), where shock effects can create large upswings but even greater downswings. In addition, the confidence intervals that arise from Theorem 5 are depicted. For the both sides, the lower bound of the estimator corresponding to high volatility is separated from the upper bound of the of the estimators, corresponding to low and mid volatility, within the stable regions. This supports the claim the the the WTI log-returns change tail index according the implied volatility. Since the conditional Hill estimator can be viewed as a function of the covariate value $x$ for a fixed sample fraction $k_{n}$ bivariately, we also plot these “risk profiles”, where slightly noisy right-skewed smiles can be seen, which indicates that the tail index is large when volatility is low and even larger when high.
@@ 7 Conclusion and outlook
Apart from considering stronger intertemporal dependence structures, our work also allows us to consider extensions in several directions. First, having an estimate of the entire tail at our disposal our results can be used to construct and statistically analyse conditional risk measures such as conditional expected shortfall or conditional tail moments. Secondly, a promising direction is to consider specific models where the same time series can be used as the covariate and response, but at different lags; Markov processes are particularly attractive in this connection. Finally, removing top observations can sometimes lead to the construction of stable estimators that allow transparent threshold selection as in Bladt et al. (2020), which is yet to be explored both for univariate and conditional cases.
@@ Appendix B Proofs of results from Section 4
using Bingham et al. (1987) Proposition 1.3.6 (i). Hence $\log\left(n/k_{n}\right)\sim(1/\gamma(x))\log(u_{n}^{x})$ and consequently
@@ Appendix E Background on fundamental limit theorems
We introduce $\alpha$- and $\beta$-mixing using Rosenblatt (1956); Volkonskii and Rozanov (1959) as primary references (see also Bradley (2005); Kulik and Soulier (2020)) and state useful results on $\beta$-mixing sequences.
@@
In this section we present existing results on regarding Central Limit Theorems using Billingsley (2009) and van der Vaart and Wellner (2023) as primary reference, see also Kulik and Soulier (2020).